# JPPO: Joint Power and Prompt Optimization for Accelerated Large Language Model Services

### v1

@@ II-B Prompt Compression
  • Transmission completeness ($\mathbf{f_{2}}$), which evaluates the integrity of information preservation during the compression process. $\mathbf{f_{2}}$ is calculated as the token retained in the compressed prompt compared to the original prompt, considering the transmission channel’s BER.
@@ III-A Problem Formulation
where $\eta$ is the Bit Error Probability (BEP) that is affected by the wireless environment and the transmit power, $P_{\text{th}}$ is the maximum allowable power consumption, $T_{\text{th}}$ represents the maximum tolerable end-to-end latency, $\mathbf{f}_{\text{th}}$ defines the minimum required fidelity. The constraints are designed to ensure practical system operation. Constraint (10a) is the energy constraint that represents the total energy budget limitation at the edge device side, introducing a critical trade-off: while higher transmission power $P_{T}$ can lead to lower BEP $\eta$ and thus improved wireless transmission fidelity $\mathbf{f_{2}}$ to enhance the overall fidelity $\mathbf{f}$, the energy constraint forces a higher compression ratio $\kappa$ to reduce both the SLM/LLM inference energy cost and wireless transmission energy consumption; however, an excessively high compression ratio can result in significant information loss from the original prompt, potentially degrading both the semantic preservation fidelity $\mathbf{f_{1}}$ and LLM service quality fidelity $\mathbf{f_{3}}$. Constraint (10b) ensures the power consumption remains within the device’s power budget, Constraint (10c) guarantees that the total service latency meets real-time requirements, Constraint (10d) maintains the quality of service by enforcing a minimum threshold on fidelity. This optimization framework allows us to find the optimal balance between compression ratio and transmission power while maintaining high-quality LLM service delivery.
@@ III-B Double DQN Solution
  • State: The state information includes the current fidelity of the transmitted message, SNR, and BEP. The state information of $n_{\rm th}$ user is captured in a 3-dimensional vector: $\left[\mathbf{f}_{n}(\eta_{n}),\gamma_{n},\right]$.
  • Action: The actions include selecting compression and power levels. The action space is denoted as $\mathcal{A}=\{\mathcal{A}_{1},\cdots,\mathcal{A}_{N}\}$, where $\mathcal{A}_{n}$ is the action of user $n$ and consists of a tuple with discrete values of compression ratio and transmission power level. The compression ratio level is discrete values range from 00 to $4$. The transmission power level is discrete values ranging from 00 to $9$, which affects BEP.
@@ IV Numerical Results
We design a customized environment with variable fidelity, SNR, and BER to simulate the wireless network-aided LLM service framework. The centralized DQN agent manages the environment, which involves selecting compression and power levels. Its goal is to balance fidelity, minimize errors, and optimize power usage. We employ the LLMLingua platform, utilizing the SLM model with GPT-Neo 125M to do prompt compression and use GPT-J 6B to generate the response [16]. The simulations are carried out on MeetingBank-transcript dataset [17]. The parameter settings are listed in Table. 1.

### v2

@@ Abstract
publicationid: pubid: This conference paper was accepted by ICC 2025.
@@ II-B Prompt Compression
  • Transmission completeness ($\mathbf{f_{2}}$), which evaluates the integrity of information preservation during the compression process. $\mathbf{f_{2}}$ is calculated as the token retained in the compressed prompt compared to the original prompt, considering the transmission channel’s Bit Error Rate (BER).
@@ III-A Problem Formulation
where $\eta$ is the BER that is affected by the wireless environment and the transmit power, $P_{\text{th}}$ is the maximum allowable power consumption, $T_{\text{th}}$ represents the maximum tolerable end-to-end latency, $\mathbf{f}_{\text{th}}$ defines the minimum required fidelity. The constraints are designed to ensure practical system operation. Constraint (10a) is the energy constraint that represents the total energy budget limitation at the edge device side, introducing a critical trade-off: while higher transmission power $P_{T}$ can lead to lower BER $\eta$ and thus improved wireless transmission fidelity $\mathbf{f_{2}}$ to enhance the overall fidelity $\mathbf{f}$, the energy constraint forces a higher compression ratio $\kappa$ to reduce both the SLM/LLM inference energy cost and wireless transmission energy consumption; however, an excessively high compression ratio can result in significant information loss from the original prompt, potentially degrading both the semantic preservation fidelity $\mathbf{f_{1}}$ and LLM service quality fidelity $\mathbf{f_{3}}$. Constraint (10b) ensures the power consumption remains within the device’s power budget, Constraint (10c) guarantees that the total service latency meets real-time requirements, Constraint (10d) maintains the quality of service by enforcing a minimum threshold on fidelity. This optimization framework allows us to find the optimal balance between compression ratio and transmission power while maintaining high-quality LLM service delivery.
@@ III-B Double DQN Solution
  • State: The state information includes the current fidelity of the transmitted message, SNR, and BER. The state information of $n_{\rm th}$ user is captured in a 3-dimensional vector: $\left[\mathbf{f}_{n}(\eta_{n}),\gamma_{n},\right]$.
  • Action: The actions include selecting compression and power levels. The action space is denoted as $\mathcal{A}=\{\mathcal{A}_{1},\cdots,\mathcal{A}_{N}\}$, where $\mathcal{A}_{n}$ is the action of user $n$ and consists of a tuple with discrete values of compression ratio and transmission power level. The compression ratio level is discrete values range from 00 to $4$. The transmission power level is discrete values ranging from 00 to $9$, which affects BER.
@@ IV Numerical Results
We design a customized environment with variable fidelity, SNR, and BER to simulate the wireless network-aided LLM service framework. The centralized DQN agent manages the environment, which involves selecting compression and power levels. Its goal is to balance fidelity, minimize errors, and optimize power usage. We employ the LLMLingua platform, utilizing the SLM model with GPT-Neo 125M to do prompt compression and use GPT-J 6B to generate the response [16]. The simulations are carried out on MeetingBank-transcript dataset [hu2023meetingbankbenchmarkdatasetmeeting]. The parameter settings are listed in Table. 1.